### v1

# Generative AI Misuse Potential in Cyber Security Education: A Case Study of a UK Degree Program

Abstract
Recent advances in generative artificial intelligence (AI), such as ChatGPT, Google Gemini, and other large language models (LLMs), pose significant challenges to upholding academic integrity in higher education. This paper investigates the susceptibility of a Master’s-level cyber security degree program at a UK Russell Group university, accredited by a leading national body, to LLM misuse. Through the application and extension of a quantitative assessment framework, we identify a high exposure to misuse, particularly in independent project- and report-based assessments. Contributing factors, including block teaching and a predominantly international cohort, are highlighted as potential amplifiers of these vulnerabilities. To address these challenges, we discuss the adoption of LLM-resistant assessments, detection tools, and the importance of fostering an ethical learning environment. These approaches aim to uphold academic standards while preparing students for the complexities of real-world cyber security.
Index Terms:
ChatGPT, Cyber Security, Higher Education
I Introduction
Large language models (LLMs) represent a significant milestone in the field of artificial intelligence (AI) and natural language processing (NLP), arising from decades of research in generative modelling. These models, exemplified by architectures such as the GPT [^1] GPT: Generative Pre-trained Transformer. family, among others, are influencing myriad fields from pharmaceuticals to finance, healthcare, information technology, translation services, and the law. Services like ChatGPT, a commercialisation of GPT models, have opened up these powerful models for public consumption. According to SimilarWeb, ChatGPT is currently the 8th most visited website in the world, ahead of Amazon, LinkedIn, TikTok and Netflix [1]. LLMs leverage self-supervised machine learning to perform language understanding and generation tasks with remarkable accuracy and versatility.
The evolution of LLMs has a deep-rooted history in generative models, beginning with early probabilistic approaches like n-grams [2] that modelled language as distributions over fixed-length sequences of words. These were succeeded by neural network-based methods such as recurrent neural networks (RNNs) and variants like long short-term memory networks (LSTMs) and gated recurrent units (GRUs). These recurrent models addressed the limitations of fixed context windows by introducing mechanisms to capture sequential dependencies across variable-length contexts. However, their efficacy was largely constrained by inefficiencies in handling very long sequences [3]. The advent of the Transformer architecture by Vaswani et al. [4] represented a paradigm shift to the problem, paving the way for modern LLMs. Unlike their predecessors, LLMs are pre-trained on diverse, large-scale corpora to learn a rich representation of language, which is subsequently fine-tuned or adapted for specific tasks. This two-phase approach has enabled LLMs to excel in tasks ranging from translation and summarisation to more complex generative tasks, including creative writing and code generation.
The recent popularisation of LLM has resulted in a significant impact within higher education. Some authors have gone as far as comparing LLM services, e.g. ChatGPT, Google’s Gemini and Anthropic’s Claude, to the advent of the calculator in mathematics departments [5]. The use of LLM servies among student populations is significant. Von Garrel and Mayer [6] presented data on the use of generative AI within a large study of German university students ($n=6311$) from 395 universities. Almost two-thirds (63.4%) of such students stated that they have used AI-based tools for their studies. The greatest reported usage of AI tools was for clarifying and explaining subject content (56.3%). Other usages were in the development of research and literature studies (with 28.6%), translations (26.6%), text analysis and generation (24.8%), and problem-solving and decision-making (22.1%). When evaluating a potential AI tool for use in their studies, the students reported that error prevention (e.g. hallucinations [^2] An hallucination refers to instances where a generative artificial intelligence model produces content that is incorrect, fabricated, or nonsensical, despite appearing plausible or convincing.), the scientificity of responses, and the ability to reason were highly regarded. Table I reproduces these preferences.
In this study, we focus specifically on the area of cyber security, which has received little attention to date in the education literature in relation to LLM misuse. Cyber security professionals operate in environments where incompetence can have severe repercussions, including data breaches, financial loss and, in the most serious of circumstances, threats to national security. The exploitation of LLMs to obtain degrees without genuine competency increases the risk of unqualified individuals entering the workforce. The integrity of assessments in cyber security programs is a key factor in ensuring that graduates possess the commensurate knowledge to perform effectively in their roles.
TABLE I: Preferences when assessing generative AI tools by university students when asked “Which aspects are/were most important to you in your assessment?” from [6].
Characteristics $N$
Error prevention during output (e.g., hallucination) 3550 56.3
Degree of scientificity (e.g., citation) 4875 77.2
Logical reasoning (e.g., answers are comprehensible) 2942 46.6
Explainability of the decision (e.g., white box vs. black box) 2176 34.5
Error detection and correction during input (e.g., grammar) 1601 25.4
Price 2555 40.5
I-A Contributions
This work principally presents a quantitative analysis of a certified UK cyber security degree program with respect to its susceptibility to LLM misuse. To this end, we apply the framework by Hickey et al. [7] for evaluating assessments from a full complement of Master’s degree modules. We develop a program exposure metric, extending the work in [7], to quantitatively measure the susceptibility of a program as a whole by weighing modules and the level of LLM misuse risk associated with their individual assessments. We discuss some of the challenges associated with program delivery, e.g. block teaching and international cohorts, which may create an environment in which misuse is incentivised. Lastly, we discuss potential solutions and their benefits and drawbacks of alternative assessment formats for addressing these challenges.
I-B Ethics
This study received ethical approval by Newcastle University on 20th April 2023. The results have been shared with the Faculty, the Degree Program Director, and the Secure and Resilient Systems (SRS) teaching group to account for potential risks, such as inadvertently influencing student assessment outcomes or exacerbating the misuse of automated tools.
II Literature Review
Much recent work has explored the misuse and potential benefits of LLMs in education. To the best of our knowledge, we are not aware of any research focussing on cyber security; however, work has been conducted in other areas of higher education. In the area of legal education, Nelson [5] argues that encouraging learners to avoid LLM abuse is to make legal education “such a joyful, meaningful, and engaging enterprise that learners will want to do it themselves – rather than let computers have all the fun.” The author identifies that LLMs are capable of passing the bar exams in the United States [8] and the Solicitors Qualifying Examination used in England and Wales [9]. Nelson leans heavily on the calculator analogy in the 1980s, stating that “it took decades for teachers to come around to the idea that using calculators in classrooms would not lead to the ‘dumbing down’ of learners, let alone embrace the idea that calculators could actually enhance mathematical education.” Indeed, the use of LLMs is considered as an opportunity for improving assessment design and student learning, rather than as a problem to be policed.
Mortlock and Lucas [10] examine the use of LLMs, specifically ChatGPT, in pharmacy education. ChatGPT was noted to return passing results for undergraduate third-year medical exams and professional pharmacy board accreditation exams in the US (see also Gilson et al. [11]). The authors raise particular concerns over: (1) the “possible hindrance of essential skill development” in pharmacy; (2) the presence of inaccuracies in LLM outputs; (3) inconsistencies between multiple versions of LLMs; and (4) the unfair advantage for students who can access the technology, given the subscription costs of such services. (See the recently released ChatGPT Pro, costing \$200 USD per month, which provides users with extra computing power for “better answers to the hardest problems problems” [12]). The authors provide recommendations and insights into potential productive uses of LLMs to bridge these concerns, such as scenario generation, simulations, and role-playing within authentic learning environments.
Perkins et al. [13] examined the extent to which academic staff could identify AI-generated content in university assessments using the Turnitin AI detection tool. Experimental submissions were created using OpenAI’s ChatGPT, employing dedicated methods to evade detection by commercial AI tools. Submissions were evaluated by 15 academic staff members alongside genuine student work. The tool flagged 91% of the experimental submissions as containing some degree of AI-generated content, but only 54.8% of the content could be actually be distinguished correctly. Academic staff referred 54.5% of these submissions to academic misconduct processes. The grades were similar between AI-generated content and student submissions. The work recommends rethinking assessment strategies to account for the rise of generative AI, such as reducing reliance on assignments susceptible to misuse.
Rogers et al. [14] recently presented a study of students’ use of ChatGPT in university-level computer science education. An online survey of $n=70$ students was conducted, exploring their current usage of ChatGPT and their ethical stance. 98.6% of students had heard of ChatGPT and that 63.8% had an active account; students disclosed that they were principally using it as a search engine and to help learn concepts that may have been unclear in class. The authors state that students were “generally…using it to learn. There is, however, a small, but noteworthy minority of students that find little moral issue with submitting work produced by the tool.” Strikingly, most students (65%) were undecided or disagreed that using ChatGPT to write assignments was a form of cheating. Furthermore, only 28% of students disagreed that using ChatGPT to write assignments was a violation of university ethical standards. Indeed, the work reports that a “notable minority of students admit to using it unscrupulously and to the potential for it to interfere with their learning.”
A significant challenge is for students to honestly disclose their true use of LLM services; it is possible that student surveys, focus groups and similar methods may suffer from social desirability bias, where participants underreport socially undesirable behaviors (like cheating) or overreport socially desirable ones due to fear of judgment or guilt, despite assurances of confidentiality [15]. This is a longstanding, problematic challenge in the evaluation of academic integrity issues more generally [16, 17, 18]. In this work, we aim to provide a data point to this end. We do not intend to measure students’ perceptions or their actual (mis)use of tools directly. Instead, we focus on a evaluating the extent to which an existing, certified program is susceptible to LLM misuse at a module and program level based on the distribution and nature of its assessments.
III Security Education as an Area of Concern
Cyber security is inherently a practice-oriented discipline, where theoretical understanding must be complemented by practical experience. Academic programs in the field typically incorporate practical components such as penetration testing labs, vulnerability assessment exercises, and incident response simulations [19, 20, 21]. Such activities are designed to cultivate an in-depth understanding of attack vectors, defensive techniques, and the real-time decision-making processes required to mitigate threats within an authentic learning environment.
The area of cyber security education, however, is relatively nascent compared to traditional fields such as philosophy or mathematics, or even computer science. It is only since the 1990s that the field has found a place in mainstream higher education; for cryptography, for instance, it was as late as approximately 2000 that the study of key techniques could be studied legally outside of the realm of national security and intelligence agencies (e.g. see Bernstein v. United States [22]). The first modules and degree programs in cyber security appeared in the mid-1990s; comparisons have been made with the accounting and finance with regards to its applied and potentially consequential nature [23].
Technical cyber security education shares similar challenges to those faced by computer science programs (e.g. see [14]). For example, students might be tempted to rely on LLMs to generate solutions or reports for practical assessments, such as scripting exploits or identifying vulnerabilities in simulated network environments. This misuse carries the risk that learners will (i) acquire superficial or incorrect knowledge, (ii) fail to develop core investigative and analytical skills, and (iii) bypass formative experiences crucial to professional competence. These factors are shared more broadly other fields of education, such as medicine [10, 11] and the law [5].
The repercussions of insufficient or deficient training are significant as cyber security professionals often assume critical roles in organisations across various sectors, including financial services, healthcare, and government. In many environments, an inability to correctly diagnose and respond to attacks can lead to severe data breaches, operational losses, or national security breaches in the most sensitive roles. A recent survey by the UK government Department for Science, Innovation and Technology (DSIT) [24] suggests that most frontline security operations and consultancy roles require a graduate degree. Approximately 42% of cyber security firms have personnel with specialist higher education degrees in the area; and 1/4 staff members possessed a specialist IT or cyber security degree. Further afield, a US study of over 11,000 job postings in 2020 estimated that approximately 60% of entry-level roles required a college degree, 24% a graduate degree, and 29% preferred a specialist cyber security certification [25].
Because professionals are often entrusted with safeguarding sensitive information and the operational security of organisations, it is critical that employers can rely on the credibility of assessments to validate a candidate’s qualifications and capabilities. The abuse of LLMs to manipulate assessment outcomes undermines this trust, making it more challenging to distinguish genuinely competent professionals from those who have improperly obtained their credentials. From a program design perspective, ensuring that students genuinely develop technical, analytical, and strategic competencies is vital; teaching staff must strike a balance between theoretical instruction and hands-on activities, while also introducing rigorous safeguards against the misuse of automated tools. If graduates lack the requisite theoretical and practical expertise due to LLM misuse, they pose a tangible risk to the organisations and infrastructures they are tasked with protecting.
TABLE II: Case study course modules. Rows give modules taught in block delivery; M10 and M11 are delivered traditionally in Sem. 2 and 3 respectively. Credit values in brackets.
Sem. 1 Sem. 2 Sem. 3
[M1] Information Security & Cryptography (10)
[M2] Risk & Trust (10)
[M7] Security of Complex Systems (10) [M10] Advanced Topics (10)
[M11] Dissertation Project (60)
[M3] Systems Security (20) [M8] Research Methods and Group Project (20)
[M4] Secure Software Development (10)
[M5] Cloud Computing (10) OR [M6] Machine Learning (10)
[M9] Network Security & Ethical Hacking (20)
IV Case Study: M.Sc. Cyber Security Program
TABLE III: Case study module descriptions.
ID Title Credits Description
[M1] Information Security & Cryptography 10 Covers fundamental techniques including cryptography, PKI, and access control. Emphasises core technical concepts and formal security methods.
[M2] Risk & Trust 10 Introduces systematic risk assessment, control measures, and security policies. Students analyse legal, ethical, and professional issues in security and trust management.
[M3] Systems Security 20 Covers system vulnerabilities, threat modeling, and security engineering. Topics include cryptographic techniques, distributed system security, and cyber-physical systems.
[M4] Secure Software Development 10 Explores secure software engineering principles such as model-driven security and formal verification. Students learn to integrate secure practices into the software development lifecycle.
[M5] Cloud Computing 10 Focuses on the principles of cloud computing, including architecture, virtualization, and security challenges. Topics include scalable computing patterns, virtual machines, and secure service design. Case studies provide practical insights into real-world applications.
[M6] Machine Learning 10 Provides applied knowledge of machine learning, including classification, regression, and clustering. Students complete a project applying machine learning to an authentic problem.
[M7] Security of Complex Systems 10 Addresses securing interconnected systems like IoT and smart grids. Topics include vulnerability analysis, human factors, and advanced attacks such as side-channel exploitation.
[M8] Research Methods and Group Project 20 Students complete a group project applying security principles to real-world problems. Develops students’ knowledge and skills in research methodologies, evaluation techniques, and collaborative project work culminating in a software artefact.
[M9] Network Security & Ethical Hacking 20 Focuses on penetration testing, ethical hacking principles, and network security. Students gain hands-on experience with vulnerability assessments and security tools.
[M10] Advanced Topics 10 Examines advanced trends like AI, Internet of Things, post-quantum security, and distributed ledger technology. The module emphasises critical evaluation of research and future trends.
[M11] Dissertation Project 60 A capstone project involving independent research under supervision. Students develop a security-based artefact, e.g., application, algorithm, user study, culminating in a dissertation.
  • In Semester 1, students select either [M5] or [M6] as an elective module, not both.
Our work considers a M.Sc. Cyber Security at a Russell Group University within the United Kingdom. The degree has been certified by a nationally leading external authority with regards to the quality and rigour of its content. The degree program has an enrollment of approximately 20–40 students per annum of which $>$85% are international students. Oftentimes, students are studying at the institution using English as the working language for the first time.
In the target program, students must study a total of 180 credits across the academic year, split into 60 credits per semester as shown in Table II. The first semester is focussed on learning core competencies in cyber security fundamentals, such as key security notions (e.g. confidentiality, integrity, availability), risk, trust, and an introduction to cryptography and system security. Students may select one 10-credit module outside of these fundamental modules on Cloud Computing or Machine Learning. The second semester involves more advanced modules, such as the security of complex systems, network security and ethical hacking, and a research methods module. Semester 2 also contains the program’s only module taught in a traditional linear mode, i.e. lasting the whole semester. The final semester is dedicated to the completion of an independent research project culminating in the production of a dissertation and demonstration.
The case study program is taught principally in block mode: a total of 8/10 modules (80%) worth 110/160 credits ($\approx$70%) are taught in this format. A small, 10-credit group-based advanced topics module and the dissertation project are the only exceptions to this delivery mode. This is not unusual within the university; the vast majority of postgraduate taught modules are delivered using block teaching within the department concerned.
IV-A Assessment Breakdown
The M.Sc. Cyber Security program is structured to develop students’ theoretical knowledge and practical skills. This is assessed through a variety of means, from written essays, traditional examinations, oral examinations, lab reports, group reports, online quizzes, and presentations. The assessments vary in type, complexity and weighting. Note that there is no strict limit on the number and type of assessments that may be used, e.g. at a School level; however, each module has at least one graded summative assessment. Many modules also contain formative assessments, which have nil contribution to students’ overall grades, but which are designed to support their learning. Our work concentrates on the summative assessments that contribute to their overall degree outcomes. We briefly describe the assessments for each module and their types in Table V.
IV-B How Risky Are The Assessments?
TABLE IV: Scoring LLM-exposed assessments from [7].
Assessment Type Location Indicator Score
Attendance In Person 0
Class Test In Person 0
Examination In Person 0
Fieldwork In Person 0
Oral Examination In Person 0
Practical Examination In Person 0
Presentation In Person 0
Seminar In Person 0
Studio Examination In Person 0
Lab Report Blended 2
Multiple Choice Questionnaire Blended 2
Portfolio Blended 2.5
Group Project Blended 3
Journal Blended 3
Assignment At Home 5
Continuous Assessment At Home 5
Essay At Home 5
Project At Home 5
In this work, we adopt the framework of Hickey et al. [7], which estimates the potential exposure of higher education assessments to LLM misuse and is intended to support the “planning and reviewing teaching and learning practices and policies”. An assessment is judged to be exposed to LLM misuse based on the following criteria:
  • The assessment type, e.g. an oral examination is deemed to have lower exposure than a take-away essay.
  • The contribution of the assessment to the module grade.
  • The intended location in which the assessment occurs. A practical examination in the presence of invigilators, for instance, is considered to have lower exposure than an assignment that is completed at home by the student where no supervision or invigilation is in place.
Hickey et al. [7] categorise the risk of various assessment types, given in Table IV. Their work gives a module exposure estimate metric for estimating the vulnerability of assessments:
Module Exposure $\text{Module Exposure Estimate}=\sum_{n=1}^{N}R_{n}S_{n}$ (1)
Where $R_{n}$ is the exposure score for a given assessment based on Table IV, $S_{n}$ is the contribution of that assessment to the final module grade as a percentage, and $N$ is the number of assessments for a given module.
Other informal measures have been suggested for evaluating assessments [26, 13, 10]. However, we adopt a quantitative approach to help easily compare modules with respect to their exposure to LLM misuse. The aforementioned exposure metric (Eq. 1) yields scores in the range $[0,5]$, where 0 represents low risk and 5 the highest risk to misuse. We applied this to all of the module assessments in the program case study, mapping the assessments to their type based on the module outline forms. We then give LLM Exposure Scores (LLMESs) on two bases: firstly, on the basis of individual assessments, from Table IV; and, secondly, the weighted LLMES for each module using Eq. 1. Table V presents the data following the application of this procedure. Histograms for both individual assessment- and module-level scores are given in Figure 1.
IV-C Findings
The results indicate that the program’s susceptibility to LLM misuse is significant. Only a single module, [M1], has a low-risk assessment—a traditional written examination—while all other modules contain some form of blended or at-home assessment. The latter types are known to be susceptible to LLM misuse due to the absence of invigilation and oversight that provide assurances that assessments were completed appropriately [26, 7]. We draw particular attention to the final project module, [M11], which is a staple of UK degree programs. The results indicate that this module is particularly vulnerable to LLM misuse. This is not surprising: the dissertation is an assessment conducted independently with limited supervision and which contributes heavily to overall module. Interestingly, the dissertation is also based around an artifact, such as a software application, which itself may be generated in whole or in part by a LLM. Other modules with report-heavy assessments, such as [M2], [M6] and [M7] also exhibited high LLM exposure scores.
Three modules were of moderate-to-low exposure, which used lab reports based on practical exercises. Only four assessments in the entire program were deemed to have low exposure ($=$0) of LLM misuse, while six assessments had the highest exposure ($=$5). Concerningly, the modal score was that assessments had very high misuse exposure potential. Some low-exposure assessments were diluted when computing module-level scores. This is because high-exposure assessments sometimes took greater precedence over low-exposure assessments, such as [M10]’s Presentation and [M11]’s Oral Examination, due to the weighting of module assessments. In general, the program contained a disproportionate number of very high-exposure assessments due to its heavy focus on take-home written assessments and project work.
TABLE V: Assessment breakdown and LLM Exposure Score (LLMES) for case study modules.
Module Type Assignment Description % LLMES (Assessment) LLMES (Module)
[M1] Written Exam Traditional examination covering relevant topics on information security and cryptography. 100 0 0
[M2] Report 1000 words, plus listings, references, and figures. 100 5 5
[M3] Quiz Canvas online quizzes on system security. 40 2 3.4
Report 1500 words with listings and documentation. (10 pages max.) 60 5
[M4] Lab Report Exercises in secure programming, logic, and verification. 100 2 2
[M5] Lab Report Testing of programming skills; code demonstration + report. 40 2 0.8
Exam Invigilated, computer-based examination. 60 0
[M6] Project Machine learning project implementation and report. 100 5 5
[M7] Assignment A 2000-word assignment on complex systems security. 100 5 5
[M8] Group Project Group literature review (20 pages max.) 30 3 3
Group Project End of module, group report (13 pages max.) 70 3
[M9] Lab Report A security analysis relating to individual exercises completed in the practical sessions. 100 2 2
[M10] Presentation Group presentation where all group members must take part. 25 0 2.25
Group Project Group report on an advanced cyber security topic with an individual reflection. 75 3
[M11] Assignment Interim report, with aim, objectives, and plan. 5 5
Oral Exam 30 min. viva (20 min. presentation + 10 min. Q&A). 25 0 4
Project Independent research project on cyber security. 75 5
Figure 1: Distributions of LLM exposure scores (LLMESs) for individual assessments (1(a)) and on a module basis (1(b)).
IV-D A Program-level Perspective
The above results focus on building assessment- and module-level exposure scores. We can also consider another metric that evaluates the susceptibility of a program in general to LLM misuse. To this end, we extend the scoring metric in Eq. 1 from [7] to weight modules’ assessment contributions to the program as a whole:
Program $\text{Program Exposure}=\sum_{i=1}^{M}\frac{\text{Credits(}i\text{)}}{\text{Total Credits}}\sum_{j=1}^{N_{i}}R_{i,j}\cdot S_{i,j}$ (2)
In our case study, for example, ‘Total Credits’ would equals 180 in line with the standard program credits for a postgraduate UK degree. ‘Credits($i$)’ represents the number of credits for the $i^{th}$ module on the program. This is a useful value to consider because the effects of high-risk modules may be reduced if their contribution to the overall program is smaller on aggregate. We computed the results for the case study using this equation. This yielded Program Exposure values of 3.10 and 3.34, depending on whether the student takes [M5] or [M6], respectively, as an optional module. As such, we can characterise the program as having ‘High’ to ‘Very High’ exposure to LLM misuse if one extends the LLMES indicators to a program level (see Table IV and Fig. 1 for bandings).
It is important to note that Hickey et al. [7] does not specifically consider other relevant variables in their scoring. For example, the work does not take into account the distribution of cohorts (e.g. home vs. international students), students’ native languages, past experiences of higher education, module delivery modes, and the effects of deadline scheduling. More broadly, they do not consider the availability of academic integrity training, quality of student support services, and access to learning resources.
It is, therefore, reasonable to hypothesise that the program’s actual exposure to LLM misuse may be significantly greater arising from the pressures induced by these additional factors. However, further research is required to understand the precise effects this may have on the tendency for students to (mis)use LLM tools.
V Potential Solutions & Recommendations
V-A Alternative Assessment Formats
One response is to adjust the assessment formats to those deemed to be at lower risk of LLM misuse. Most straightforwardly, this may involve a greater utilisation of traditional written examinations, oral examinations, seminars and in-class tests. These assessments can be more strictly invigilated within environments where LLMs can be not be deployed and misused. Informal conversations suggest that this was being planned in some module; however, it is important to emphasise the deleterious effects that short-term examinations can have on student and instructor well-being [27, 28].
We believe the use of block teaching within the case study has relevance with respect to the potential misuse of LLMs. It has been reported that intensive teaching modes, i.e. over 1–4 week periods, hinders active learning, and places students under additional fatigue and stress compared to traditional forms of delivery in higher education [29, 30, 31]. We can infer that students may resort to LLMs as an attractive tool to alleviate these pressures, creating an environment in which misuse may occur. The program is also taken by a large cohort of international students, who may be writing assessments in English at a UK university for the first time, which is an additional factor. We hypothesise that these constraints significantly increase the likelihood of LLMs being misused by students at the time of assessment. However, this warrants a dedicated study, which we defer to future research.
Staff may wish to consider implementing alternative assessment formats that are more resistant to automation. These could involve designing tasks that mirror real-world challenges and scenarios, equipping students to navigate the complexities and uncertainties they are likely to face in their professional roles. Examples include performance-based assessments such as simulations, case studies, and practical skill evaluations, requiring students to demonstrate their competencies in applied contexts. Similar approaches have been advocated in fields like pharmacy and medicine, where hands-on expertise is crucial [10, 11]. The cyber security education literature is rich with relevant examples; for instance, the use of ‘capture the flag’ (CTF) [32], wargaming [33], and red/blue team simulations [34], which immerse learners in a manner closely aligned with the complexities of real-world cyber environments. However, it is important to note that such assessments are generally complex to design, can lack scalability to large class sizes, and impose significant additional IT costs.
V-B Detection Methods
Another approach to mitigating the misuse of LLMs involves deploying detection mechanisms capable of flagging AI-generated content. Several commercial and open-source tools exist that claim to detect passages produced by language models through statistical analysis of text features or the identification of linguistic patterns unique to automated systems. Although these tools offer an appealing method for preserving academic integrity, their efficacy and reliability remain quesitonable: recent evaluations suggest that detection algorithms are prone to both false positives and false negatives, particularly as LLMs evolve and students become aware of more sophisticated evasion techniques [35, 36, 37].
Consideration has also been given to tailoring assessments such that LLMs would generate an easily identifiable incorrect response. Such responses would then provide evidence that LLM misuse occurred, effectively ‘booby trapping’ assessment specifications that students have entered verbatim into LLM services. However, the ethics and pedagogical value of such approaches are highly contentious. If students are unaware that certain assessment elements are effectively designed to deceive or trick LLMs, then staff may foster an adversarial dynamic that undermines student trust. While assessment ‘booby trapping’ is academically intriguing, its adoption requires further research particularly on ethical and pedagogical aspects.
Moreover, an focussing on detecting misuse overzealously may undermine student trust. Students may feel unfairly targeted by automated screening methods, especially in diverse international cohorts where language proficiency varies, and where LLMs may be used legitimately (such as for proofreading). Furthermore, any detection system—however sophisticated—is subject to non-negligible error rates. The resulting interventions may be contentious if students are penalised based on an uncertain evidence base.
V-C Limitations
The results presented in this study are subject to some limitations. Reducing programs to a single value of LLM exposure, while useful for comparative purposes, particularly for university policymakers and faculty decision-makers, may oversimplify the complexities inherent in assessment design and delivery. Such a metric may obscure variations in assessment quality; a module with a high LLM exposure score might still produce good learning outcomes if it is supported, for example, by a well-designed rubric that mitigates the limitations of current LLMs, such as their propensity for vague responses and misleading citations.
The approach of weighing programme susceptibility based on credit allocations represents a first attempt at quantifying LLM-exposed programs, but it introduces its own challenges. It assumes that modules’ credit values are equivalent to their contribution to students’ overall learning. Modules with lower credit weights but important practical elements may be disproportionately undervalued, while heavily weighted theoretical modules could skew the overall program exposure metric. The metric also assumes a static model of LLM capabilities and does not account for their evolution. LLMs have already, for example, shown promise in crafting realistic technical presentations and generating real-time verbal responses to questions (e.g. as in a remote oral examination). Ongoing refinement and reflection is thus important in ensuring responsible use of the metrics and results given in this work.
VI Conclusion
The rapid proliferation of LLMs presents both challenges and opportunities in cyber security education. We focussed on the risk to academic integrity, presenting a case study of a certified UK Master’s program whose susceptibility to LLM misuse was examined using a quantitative approach. We introduced a new program-level metric that reveals a high degree of susceptibility to LLM misuse, a pattern likely echoed in other UK programs. This is due principally to the heavy use of traditional assessment modalities, such as take-home reports and project-based work, increasing the aggregated program-level risk. The analysis was agnostic to contextual factors, such as the block teaching formats and highly diverse student cohorts. These may create an environment where students see LLMs as tools for alleviating academic stress, which we consider as important future research.
Addressing these problems is not straightforward. It is prudent that programs encourage assessment formats that are resistant to LLM misuse. In cyber security, this may involve favoring authentic, performance-based tasks such as red/blue team simulations. Nevertheless, these can require dedicated expertise and IT resources to work well, which must be balanced against the pragmatic benefits of traditional assessments. More traditional methods, such as oral and written examination, could also be considered as part of an holistic solution in reducing LLM exposure risks, but their drawbacks must be recognised by module and program designers. Furthermore, while detection and monitoring strategies can be helpful, they should be employed responsibly in order to uphold trust in the learning experience. It is important for institutions to raise the awareness of ethical LLM usage, providing structural support to mitigate the pressures driving misuse.
Acknowledgment
Carlton Shepherd thanks the Faculty of Science, Agriculture and Engineering (SAgE) at Newcastle University for funding this work though an internal education grant. He also thanks Essam Ghadafi and other members of the Secure and Resilient Systems (SRS) group within the School of Computing for insightful discussions on the topic.